OpenStudy (anonymous):

A 20-Kg rock is at the edge of a 100 m height cliff. a) What is the rock kinetic energy respect to the base of cliff? b) if the rock is falling. What is its kinetyc energy right before crashing against the ground?

OpenStudy (anonymous):

a) 0 J b) 20000 J

OpenStudy (anonymous):

can you show the work, please?

OpenStudy (anonymous):

im really confused

OpenStudy (anonymous):

see the rock is stationary on the cliff..so its kinetic energy is zero...get it?

OpenStudy (anonymous):

you mean that no movement = kinetic energy =0?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what about the second one?

OpenStudy (anonymous):

see now when the rock falls, the only force acting on it is gravity, which is a CONSERVATIVE force...so the total mechanical energy remains conserved at all points

OpenStudy (anonymous):

in that case ME (initial) = ME(final)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

KE(i) + PE(i) = KE(f) +PE(f) 0 + mgh = KE +0 KE = 20(10)(100)

OpenStudy (anonymous):

get it?

OpenStudy (anonymous):

what does the "i" and "f" inside the brackets mean?

OpenStudy (anonymous):

initial and final

OpenStudy (anonymous):

does this equation works for all cases when something falls?

OpenStudy (anonymous):

yes..coz only gravity acts..which is a conservative force

OpenStudy (anonymous):

ill tell u one more concept which can be used....anywhere

OpenStudy (anonymous):

should i?

OpenStudy (anonymous):

sure, thank you so much. I really need that.

OpenStudy (anonymous):

its called the work-energy theorem The total work done by all the forces acting on a body equals the change in its kinetic energy

OpenStudy (anonymous):

try doing this with the nw concept

OpenStudy (anonymous):

I have to solve tree exercises. Can you help me with them first, please?

OpenStudy (anonymous):

wts tree exercise?

OpenStudy (anonymous):

three exercises. Dam Ecuadorian that cannot speak English very well lol

OpenStudy (anonymous):

can you help me?

OpenStudy (anonymous):

no prob as long as the battery on my laptop allows...fire away quickly

OpenStudy (anonymous):

Initially, a 2 Kg rock is at reast. It losses 400J of potential energy as it falls to the ground. a) Find the kinetic energy that the rock gains as it falls b)what is the speed of the rock right before it hits the ground

OpenStudy (anonymous):

since energy is again conserved Loss in Potential Energy = Gain in kinetic energy

OpenStudy (anonymous):

400=KE=0.5mv^2 400=v^2 v=20

OpenStudy (anonymous):

get it?

OpenStudy (anonymous):

No 100% but today im not sleeping in order to undertand it. Help me with the last one

OpenStudy (anonymous):

go on

OpenStudy (anonymous):

Mary weights 420N and is sitting on a swing . The swing is 0.4 m from the ground. Thomas push the swing until it is at 1m from the ground.

OpenStudy (anonymous):

a) what is the speed of the swing (mary) when it reaches the lowest position?

OpenStudy (anonymous):

mg(1-0.4) = 0.5mv^2 g(0.6) = 0.5v^2 v=root(12)

OpenStudy (anonymous):

b) if in its lowest position the speed of mary is 2m/s.. What work does do friction do over the swing?

OpenStudy (anonymous):

now use Work-Energy Theorem W(friction) + W(gravity) = Change in KE W(friction) + mg(0.6) = 0.5 x m x 2^2 W(friction) = 2m -6m =-4m = -168 J

OpenStudy (anonymous):

this theorem is very useful

OpenStudy (anonymous):

Thank you very much. Im gonna review the math behind work and energy and then im going to try to understand the problems you helped me with. Thanks him1618 im becoming you fan :)

OpenStudy (anonymous):

its very easy when u start playing around with calculus

OpenStudy (anonymous):

calculus?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

understanding math in work and energy and the whole of mechanics is easy when u play with calculus

OpenStudy (anonymous):

Im not so bad at calculus; but i dont understant that lol

OpenStudy (anonymous):

for eg

OpenStudy (anonymous):

dW = Fdx = mdvdx/dt (as F = mdv/dt) = mvdv(as dx/dt=v) = 0.5mv^2

OpenStudy (anonymous):

so u c that the work done by all forces is the change in kinetic energy

OpenStudy (anonymous):

ahhh got you. Do you think I can handle potential and kinetic energy in one week? Is that easy?

OpenStudy (anonymous):

My maths are no so bad. Physic topics are more difficult for me to understand

OpenStudy (anonymous):

physics and maths correlate

OpenStudy (anonymous):

just do it topicwise

OpenStudy (anonymous):

Work and its definitions Conservative and NonConservative Forces Potential Energy Conservation of Energy Work-Energy Theorem Power

OpenStudy (anonymous):

Ok. By the way. Where are you from? People always help me and I dont even know where they are from . Im Luis and Im from Ecuador :)

OpenStudy (anonymous):

Himanish from India

OpenStudy (anonymous):

Nice. country! So I really appreciate you help Himanish. Im more relaxed now because of your help. Now im gonna sleep coz it is 2 am here. Hope we get in touch another day. Thank you

OpenStudy (anonymous):

welcome...12 pm here..gnite though

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

hey

OpenStudy (anonymous):

Hi Himanish!

OpenStudy (anonymous):

what time is in your country now. Here is almost 2 PM.

OpenStudy (anonymous):

11.45 pm

OpenStudy (anonymous):

are goint to sleep now. Or you work like me, at nights?

OpenStudy (anonymous):

night

OpenStudy (anonymous):

haha. That is cool dude. I cannot concentrate in the day

OpenStudy (anonymous):

me too

OpenStudy (anonymous):

so u got thru planning for work and energy?

OpenStudy (anonymous):

Kind of. I just started reading that. I had to help my mom in the morning. Now Im free. I was reading about maximun and minima of a function in the morning one hour ago...

OpenStudy (anonymous):

are you teacher? or are you student?

OpenStudy (anonymous):

student

OpenStudy (anonymous):

hey him. can you help me with this exercise? The radius of the earth is 6370Km and the radius of mars is 3440Km. If an object weights 200N on earth, what will be the gravity on mars? Mars has a mass of 0.11 times the mass of the earth.